A contractor, who has $12 million dollars budget, is planning for his 2019 investments. The average after tax yearly profit for public school and residential construction is 7% and 8%, respectively. The third option is to invest in the stock market where the average yearly profit is predicted to be 12% (after tax). In order to benefit from tax exemptions, the tax advisor recommends that contractor’s involvement in public school construction be at least 3 times as compared to residential building. Investment in stock market is very risky, so the contractor is not willing to put more than $ 2 million in it. What is the best 2019 investment scenario? Calculate the corresponding profit. (use graphical procedure)
Answer #1

Amount invested in Stock market is $2m which will have a return of 12%, ie $240,000

For the balance $10m, amount to be invested in in school construction should be $10m/4 X 3 = $7.5m, which will have a return of 7%, ie $525,000

Remaining amount of $2.5m will be invested in residential construction which will have a return of 8%, ie $200,000

Total return is 240,000+525,000+200,000= $965,000

Rate of return is 965,000/12,000,000 = 8.04%